I need help with a problem for logarithms.
x-1/x+2 = 3/4

Answer by dabanfield(803)

I need help with a problem for logarithms.
x-1/x+2 = 3/4
Multiply both sides of the equation aboove by x+2:
((x-1)/(x+2))*(x+2) = (3/4)*(x+2)
The (x-2) factor on the left cancels:
x-1 = (3*(x-2))/4
Multiply both sides above by 4:
4*(x-1) = 3*(x-2)
4x - 4 = 3x - 6
Solve for x. No need to use logarithms.